OpenStudy (anonymous):

What is 3X-Y=13 in Slope Intercept form?

OpenStudy (mathstudent55):

You need an equal sign somewhere. Is it 3x - y + 13 = 0 or 3x - y = 13? Either way, to put this equation in slope-intercept form, solve for y.

OpenStudy (anonymous):

Oops, I meant 3x-y=13. Sorry about that! What I got was y=3x+13. Is that right?